# Sampling as Bandits: Evaluation-Efficient Design for Black-Box Densities

### v1

@@ B.6 Proof of Corollary˜3
As shown in the proof of Corollary˜2, the MMD of the kernel $\kappa$ metrises the weak convergence. With no loss of generality, suppose that $\epsilon$ is small enough that $\sqrt{\log(2/\epsilon)}>1$. Let $C$ be a constant larger than $C_{1}$, $C_{2}$, and $C_{3}M$. Since $N\geq 1$ and $M\geq 1$, we have
$d(p,\delta_{N})\leq\frac{C_{1}+C_{2}\sqrt{\log(2/\epsilon)}}{\sqrt{N+M}}+C_{3}\frac{M}{N+M}\leq\frac{C\sqrt{\log(2/\epsilon)}}{\sqrt{N+M}}=:\delta$ (99)
@@
$\mathbb{P}\left(d(p,\delta_{N})>\delta\right)$ $\leq 2\exp\left(-\frac{\delta^{2}}{2C^{2}}(N+M)\right).$ (100)
It then follows from the Borel-Cantelli lemma that $d(p,\delta_{N})$ converges to zero almost surely in the limit $N\to 0$, meaning that $\delta_{N}$ weakly converges to $p$ almost surely. ∎
@@ Appendix C Simulation Studies
This section presents additional simulation studies on BIS. Section˜C.1 discusses the importance of the point-selection procedure of BIS in comparison with the randomised BO. Section˜C.2 demonstrates the performance of BIS for different choices of the function $\phi$. Section˜C.3 illustrates the performance of BIS for different choices of the size of the proposal sequence $M$.
C.1 Importance of Discrete Optimisation with No Revisit
@@
In BO, the point at each iteration is selected by the optimisation over an infinite set, such as a compact subset of $\mathbb{R}^{d}$. In contrast, BIS select the point by the optimisation over a discrete set under the constraint that points selected in the past will never be revisited. The optimisation on the discrete set, together with the constraint, assists in efficient exploration of the domain of interest. It prevents cases where the point to be selected next will be arbitrarily close to the points already selected in the past. In the long run, therefore, the points selected in BIS do not concentrate on a particular subset of the domain. This is vital for the convergence guarantee of the importance sampling scheme.
@@
The randomisation approach proposed in Kim and Sanz-Alonso (2024) extends BO with the aims of approximating posterior densities. They proposed performing BO whlie adding an extra random point in the training dataset of GP at each iteration. This means that at each iteration it obtains two points, where one of then is the point acquired by BO and the other is a random point. They showed the convergence of the GP posterior mean to a target posterior, when the GP is well specified. However, we shall demonstrate that importance sampling based on the points selected by such a randomised BO fails to recover the target posterior.
@@
Figure˜7 visualises 100 points obtained by the randomised BO of Kim and Sanz-Alonso (2024), for each benchmark density in Section˜5.1. At each iteration of the randomised BO, a point from the scaled Halton sequence is used as the random point to add. Following the experimental setting in Section˜5.1, the first 10 points were randomly selected from the Halton sequence. It illustrates that the points selected by BO (triangular) gradually concentrate to the maxima of the posterior density, while the random points from the Halton sequence (square) are dispersed over the domain. Figure˜8 further confirms that importance sampling based on these samples fails to converge to the target posterior. For each target density, the approximation error of the resulting weighted samples deteriorated after a certain number of iterations.
@@
The failure in the convergence roots in that points selected by BO concentrates at the mode of the target posterior after some iterations. Thus, after some iterations, the majority of the samples used for importance sampling are enforced to have one particular value, which is the mode. Moreover, those majority points sharing the same value are assigned the highest importance weights, since the mode is the point where the posterior probability is maximal by definition. Therefore, the influence of those majority points concentrating at the mode will be dominant. Roughly speaking, this means that, after some iterations, the weighted samples in the randomised BO starts approximating the Dirac distribution at the mode rather than the target density. On the other hand, BIS, by design, prevents such a concentration of samples.
@@
Figure 7: Visualisation of 100 samples obtained by the randomised BO of Kim and Sanz-Alonso (2024) for each benchmark density. The initial 10 points are star-shaped. The following points obtained by BO and QMC are illustrated in, respectively, triangle- and square-shaped. The contour values of each density were powered to $1/3$ for better visualisation of the tail geometry.
@@
Figure 8: The approximation error of BIS (solid line) and standard importance sampling based on the points of the randomised BO (dashed line). The experiment was repeated 10 times, where the bold line represents the averaged error and the band represents the 95% confidence interval.
@@ C.2 Different Choices of Function $\phi$ in GP-UJB
In Section˜3.4, Table˜1 listed examples of the function $\phi$ used in GP-UJB. This subsection compares the performance of BIS for different choices of the function $\phi$ through a simulation study. We used the experimental setting in Section˜5.1, while changing the choice of the function $\phi$. Figure˜9 visualises 100 samples of BIS obtained for the banana density in Section˜5.1, under three different choices of the function $\phi$. It demonstrates that, under the quadratic function $\phi(x)=x^{2}$ and the relu function $\phi(x)=\max(x,0)$, a relatively large portion of samples were situated outside the high probability region of the target density. On the other hand, the exponential function $\phi(x)=\exp(x)$ efficiently situated the majority of samples around the high probability region.
Figure˜10 shows the approximation error of BIS for each benchmark density in Section˜5.1, under three different choices of the function $\phi$. While all three choices of the function $\phi$ provide a reasonable approximation error, the default choice $\phi(x)=\exp(x)$ achieved a relatively better approximation error overall. The approximation errors for the Gaussian and bimodal densities were similar, and those for the banana density exhibit a difference, as also exemplified in Figure˜9. Given this simulation study, we recommend the exponential function $\phi(x)=\exp(x)$ as the default choice to use in GP-UJB.
@@
Figure 9: Visualisation of 100 samples of the banana density obtained by BIS for different $\phi$. The contour values of the density were powered to $1/3$ for better visualisation of the tail geometry.
@@
Figure 10: The approximation error of BIS for each benchmark density when the function $\phi(x)$ in GP-UJB is quadratic $x^{2}$ (dashed line), relu $\max(x,0)$ (dotted line), and exponential $\exp(x)$ (solid line). The experiment was repeated 10 times, where the bold line represents the averaged error and the band represents the 95% confidence interval.
@@ C.3 Sensitivity of Proposal-Sequence Size $M$
The size of the proposal sequence $M$ is a hyperparameter of BIS. The convergence rates in Section˜4 imply the trade-off between small and large values of the size $M$. If the size $M$ is large, the first term in the convergence rates—regardless of the deterministic or probabilistic one—is suppressed, while the second term amplifies. Conversely, if the size $M$ is small, the first term is not suppressed, while the second term does not amplify. In the extreme case where $M$ is equal to one, BIS is identical to standard self-normalised importance sampling by construction. Choosing a reasonable value of the size $M$ controls the trade-off, leading to an efficient approximation performance of BIS.
This subsection performs a simulation study on sensitivity of BIS to the size of the proposal sequence $M$. We used the experimental setting in Section˜5.1, while changing the size of the proposal sequence $M$. Figure˜11 shows the approximation error of BIS for each benchmark density, under different values of the proposal-sequence size $M$. As anticipated, when the size $M$ was the small value $2$, 88, and $32$, the approximation performance of BIS got similar to that of standard importance sampling based on the scaled Halton sequence. In addition, the excessively large value of the proposal-sequence size $M=32768$ didn’t result in the best approximation performance of BIS. Overall, the size $M=2048$ and $M=8196$ led to the fastest decay of the approximation error of BIS. As the smaller size $M$ reduces the computation cost in BIS, we recommend the size $M=2048$ for experiments in two-dimensional domains based on this simulation study.
@@
Figure 11: The approximation error of BIS for each benchmark density when the sice of the proposal sequence is $2$, 88, $32$, $128$, $512$, $2048$, $8196$, and $32768$. The experiment was repeated 10 times, where the bold line represents the averaged error and the band represents the 95% confidence interval.
@@ Appendix D Additional Experimental Details
This section provides additional details of the experiments presented in the main text. Section˜D.1 contains that of Section˜5.1 and Section˜D.2 contains that of Section˜5.2.
D.1 Additional Details of Section˜5.1
All the benchmark densites are given in the form proposed in Järvenpää et al. (2021) s.t.
$p(\theta)\propto\exp\bigg{(}-\frac{1}{2}\begin{bmatrix}T_{1}(\theta)\\ T_{2}(\theta)\end{bmatrix}^{\mathrm{T}}\begin{bmatrix}1&\rho\\ \rho&1\end{bmatrix}\begin{bmatrix}T_{1}(\theta)\\ T_{2}(\theta)\end{bmatrix}\bigg{)}$. (101)
@@
$k(\theta,\theta^{\prime})=\sigma^{2}\exp\left(-\frac{\|\theta-\theta^{\prime}\|^{2}}{2l^{2}}\right)$ (102)
@@
$\operatorname{MMD}(p,q)^{2}=\mathbb{E}_{\theta,\theta^{\prime}\sim p}[\kappa(\theta,\theta^{\prime})]-2\mathbb{E}_{\theta\sim p,\theta^{\prime}\sim q}[\kappa(\theta,\theta^{\prime})]+\mathbb{E}_{\theta,\theta^{\prime}\sim q}[\kappa(\theta,\theta^{\prime})].$ (103)
This alternative form of MMD can be efficiently estimated by Monte Carlo integration using samples from the densities $p$ and $q$. Throughout the experiments, we used MMD associated with the Gaussian kernel $\kappa(\theta,\theta^{\prime})=\exp(-0.5\|\theta-\theta^{\prime}\|^{2}/h)$ with the constant $h=0.1$.
@@
$\operatorname{TVD}(p,q):=\frac{1}{2}\int_{\Theta}\left|p(\theta)-q(\theta)\right|d\theta.$ (104)
We approximate the integral over $\Theta$ by numerical integration. In this experiment, the densities $p$ and $q$ to be plugged in TVD are available only up to the normalising constants. We also approximate their normalising constants by numerical integration. Let $\{\eta_{i}\}_{i=1}^{N}$ be the $N$ points from the scaled Halton sequence on the domain $\Theta$. Denote the volume of the domain $\Theta$ by $V$. Let $\tilde{p}$ and $\tilde{q}$ be the proportional terms of the densities $p$ and $q$. Then their normalising constants, denoted by $Z_{p}$ and $Z_{q}$, are approximated as follows:
$Z_{p}\approx\frac{V}{N}\sum_{i=1}^{N}\tilde{p}(\eta_{i})=:\hat{Z}_{p}\quad\text{and}\quad Z_{q}\approx\frac{V}{N}\sum_{i=1}^{N}\tilde{q}(\eta_{i})=:\hat{Z}_{q}.$ (105)
@@
$\operatorname{TVD}(p,q)$ $\approx\frac{V}{2N}\sum_{i=1}^{N}\left|\frac{\tilde{p}(\eta_{i})}{\hat{Z}_{p}}-\frac{\tilde{q}(\eta_{i})}{\hat{Z}_{q}}\right|.$ (106)
@@
Table˜2 summarised the number of samples required for standard importance sampling to surpass the approximation error of 100 weighed samples obtained by BIS. For standard importance sampling, we used the scaled Halton sequence as samples and assigned self-normalised importance weights to them. It demonstrates that standard importance sampling required nearly 2,000 samples on average to achieve the approximation error that BIS does with 100 samples.
Table 2: The number of samples required for standard importance sampling to surpass the approximation error of 100 BIS weighed samples. The experiment was repeated 10 times, where the averaged number was reported together with the standard deviation in a bracket.
@@
D.2 Additional Details of Section˜5.2
The synthetic likelihood of the modified Lorentz weather model is computed using summary statistics of sample paths simulated from the model. Denote by $s\in\mathbb{R}^{d}$ summary statistics of the observed sample path, for some dimension $d$. Denote by $s(\theta)\in\mathbb{R}^{d}$ summary statistics of a sample path simulated from the model under a parameter $\theta$, where $s(\theta)$ is a random variable induced by simulation from the model. Let $m(\theta)\in\mathbb{R}^{d}$ and $\Sigma(\theta)\in\mathbb{R}^{d\times d}$ be the mean and covariance of $s(\theta)$. The synthetic likelihood at $\theta$ is then the Gaussian density $N(s\mid m(\theta),\Sigma(\theta))$ evaluated at the summary statistics $s$ of data. In practice, the mean and covariance are estimated from multiple i.i.d. realisations of the summary statistics $s(\theta)$. Thus, the synthetic likelihood is computed using multiple sample paths simulated from the model at each $\theta$. For the modified Lorentz weather model, Hakkarainen et al. (2012) used the following six summary statistics: (i) mean of $x_{k}(t)$ over time, (ii) variance of $x_{k}(t)$ over time, (iii) autocovariance of $x_{k}(t)$ with time lag one, (iv) covariance of $x_{k}(t)$ and $x_{k+1}(t)$ over time, (v) cross covariance of $x_{k}(t)$ and $x_{k-1}(t)$ with time lag one, (vi) cross covariance of $x_{k}(t)$ and $x_{k+1}(t)$ with time lag one. The summary statistics computed for each variable $x_{k}(t)$ is averaged over $k$. Thus, we have six-dimensional summary statistics for each sample path of the 40 variables.

### v2

@@ B.5 Proof of Corollary˜3
As shown in the proof of Corollary˜2, the MMD of the kernel $\kappa$ metrizes the weak convergence. With no loss of generality, suppose that $\epsilon$ is small enough that $\sqrt{\log(2/\epsilon)}>1$. Let $C$ be a constant larger than $C_{1}$, $C_{2}$, and $C_{3}M$. Since $N\geq 1$ and $M\geq 1$, we have
$d(p,\delta_{N})\leq\frac{C_{1}+C_{2}\sqrt{\log(2/\epsilon)}}{\sqrt{N+M}}+C_{3}\frac{M}{N+M}\leq\frac{C\sqrt{\log(2/\epsilon)}}{\sqrt{N+M}}=:\delta$ (S72)
@@
$\mathbb{P}\left(d(p,\delta_{N})>\delta\right)$ $\leq 2\exp\left(-\frac{\delta^{2}}{2C^{2}}(N+M)\right).$ (S73)
It then follows from the Borel-Cantelli lemma that $d(p,\delta_{N})$ converges to zero almost surely in the limit $N\to 0$, meaning that $\delta_{N}$ weakly converges to $p$ almost surely. The consistency of the Monte Carlo estimator immediately follows from the definition of the weak convergence. Here, since the domain $\Theta$ is compact, the continuous integrand $f$ is bounded by extreme value theorem. ∎
B.6 Proof of Proposition˜1
Proof.
$(\star)$ $=\int_{\Theta}\Big|\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[\phi\left(f(\theta)\right)\right]-\phi\left(g(\theta)\right)\Big|^{2}d\theta=\int_{\Theta}\Big|\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[\phi\left(f(\theta)\right)-\phi\left(g(\theta)\right)\right]\Big|^{2}d\theta.$ (S74)
It then follows from the Jensen’s inequality that
$(\star)$ $\leq\int_{\Theta}\bigg(\underbrace{\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\Big[\left|\phi\left(f(\theta)\right)-\phi\left(g(\theta)\right)\right|\Big]}_{=:(*)}\bigg)^{2}d\theta.$ (S75)
$|\phi(a)-\phi(b)|\leq\bigg(\sup_{c\in[a,b]}|\phi^{\prime}(c)|\bigg)|b-a|\leq\max\left(|\phi^{\prime}(a)|,|\phi^{\prime}(b)|\right)|b-a|$ (S76)
$(*)$ $\leq\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\Big[\max\left(|\phi^{\prime}(f(\theta))|,|\phi^{\prime}(g(\theta))|\right)\left|f(\theta)-g(\theta)\right|\Big]$ (S77)
$\leq\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\Big[\underbrace{\max\left(\exp(f(\theta)),\exp(g(\theta))\right)}_{=:(**)}\left|f(\theta)-g(\theta)\right|\Big]$ (S78)
$(*)\leq\left(\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[(**)^{2}\right]\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[\left|f(\theta)-g(\theta)\right|^{2}\right]\right)^{\frac{1}{2}}.$ (S79)
Observe the following trivial inequality
$(**)^{2}=\max\left(\exp(2f(\theta)),\exp(2g(\theta))\right)\leq\exp(2f(\theta))+\exp(2g(\theta)).$ (S80)
$\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[(**)^{2}\right]$ $\leq\exp(2m_{n}(\theta)+2k_{n}(\theta,\theta))+\exp(2g(\theta))$ (S81)
$\leq\underbrace{\sup_{\theta\in\Theta}\exp(m_{n}(\theta)+k_{n}(\theta,\theta))^{2}+\sup_{\theta\in\Theta}\exp(g(\theta))^{2}}_{=:C}.$ (S82)
Combining the bounds (S79) and (S82), we have the overall bound of the original term $(\star)$:
$(\star)\leq C\int_{\Theta}\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[\left|f(\theta)-g(\theta)\right|^{2}\right]d\theta=C\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[\int_{\Theta}\left|f(\theta)-g(\theta)\right|^{2}d\theta\right]$ (S83)
@@ Appendix C Simulation Studies
This section presents additional simulation studies on BIS. Section˜C.1 discusses the importance of the point-selection procedure of BIS in comparison with the randomized BO. Section˜C.2 demonstrates the performance of BIS for different choices of the function $\phi$. Section˜C.3 illustrates the performance of BIS for different choices of the size of the proposal sequence $M$.
C.1 Importance of Discrete Optimization with No Revisit
@@
In BO, the point at each iteration is selected by the optimization over an infinite set, such as a compact subset of $\mathbb{R}^{d}$. In contrast, BIS select the point by the optimization over a discrete set under the constraint that points selected in the past will never be revisited. The optimization on the discrete set, together with the constraint, assists in efficient exploration of the domain of interest. It prevents cases where the point to be selected next will be arbitrarily close to the points already selected in the past. In the long run, therefore, the points selected in BIS do not concentrate on a particular subset of the domain. This is vital for the convergence guarantee of the importance sampling scheme.
@@
The randomization approach proposed in Kim2024 extends BO with the aims of approximating posterior densities. They proposed performing BO whlie adding an extra random point in the training dataset of GP at each iteration. This means that at each iteration it obtains two points, where one of them is the point acquired by BO and the other is a random point. They showed the convergence of the GP posterior mean to a target posterior, when the GP is well specified. However, we shall demonstrate that importance sampling based on the points selected by such a randomized BO fails to recover the target posterior.
@@
Figure˜S1 visualizes 100 points obtained by the randomized BO of Kim2024, for each benchmark density in Section˜6.1. At each iteration of the randomized BO, a point from the scaled Halton sequence is used as the random point to add. Following the experimental setting in Section˜6.1, the first 10 points were randomly selected from the Halton sequence. It illustrates that the points selected by BO (triangular) gradually concentrate to the maxima of the posterior density, while the random points from the Halton sequence (square) are dispersed over the domain. Figure˜S2 further confirms that importance sampling based on these samples fails to converge to the target posterior. For each target density, the approximation error of the resulting weighted samples deteriorated after a certain number of iterations.
@@
The failure in the convergence roots in that points selected by BO concentrates at the mode of the target posterior after some iterations. Thus, after some iterations, the majority of the samples used for importance sampling are enforced to have one particular value, which is the mode. Moreover, those majority points sharing the same value are assigned the highest importance weights, since the mode is the point where the posterior probability is maximal by definition. Therefore, the influence of those majority points concentrating at the mode will be dominant. Roughly speaking, this means that, after some iterations, the weighted samples in the randomized BO starts approximating the Dirac distribution at the mode rather than the target density. On the other hand, BIS, by design, prevents such a concentration of samples.
@@
Figure S1: Visualization of 100 samples obtained by the randomized BO of Kim2024 for each benchmark density. The initial 10 points are star-shaped. The following points obtained by BO and QMC are illustrated in, respectively, triangle- and square-shaped. The contour values of each density were powered to $1/3$ for better visualization of the tail geometry.
@@
Figure S2: The approximation error of BIS (solid line) and standard importance sampling based on the points of the randomized BO (dashed line). The experiment was repeated 10 times, where the bold line represents the averaged error and the band represents the 95% confidence interval.
@@ C.2 Different Choices of Function $\phi$ in GP-UJB
In Section˜5.3, Table˜1 listed examples of the function $\phi$ used in GP-UJB. This subsection compares the performance of BIS for different choices of the function $\phi$ through a simulation study. We used the experimental setting in Section˜6.1, while changing the choice of the function $\phi$. Figure˜S3 visualizes 100 samples of BIS obtained for the banana density in Section˜6.1, under three different choices of the function $\phi$. It demonstrates that, under the quadratic function $\phi(x)=x^{2}$ and the relu function $\phi(x)=\max(x,0)$, a relatively large portion of samples were situated outside the high probability region of the target density. On the other hand, the exponential function $\phi(x)=\exp(x)$ efficiently situated the majority of samples around the high probability region.
Figure˜S4 shows the approximation error of BIS for each benchmark density in Section˜6.1, under three different choices of the function $\phi$. While all three choices of the function $\phi$ provide a reasonable approximation error, the default choice $\phi(x)=\exp(x)$ achieved a relatively better approximation error overall. The approximation errors for the Gaussian and bimodal densities were similar, and those for the banana density exhibit a difference, as also exemplified in Figure˜S3. Given this simulation study, we recommend the exponential function $\phi(x)=\exp(x)$ as the default choice to use in GP-UJB.
@@
Figure S3: Visualization of 100 samples of the banana density obtained by BIS for different $\phi$. The contour values of the density were powered to $1/3$ for better visualization of the tail geometry.
@@
Figure S4: The approximation error of BIS for each benchmark density when the function $\phi(x)$ in GP-UJB is quadratic $x^{2}$ (dashed line), relu $\max(x,0)$ (dotted line), and exponential $\exp(x)$ (solid line). The experiment was repeated 10 times, where the bold line represents the averaged error and the band represents the 95% confidence interval.
@@ C.3 Sensitivity of Proposal-Sequence Size $M$
The size of the proposal sequence $M$ is a hyperparameter of BIS. The convergence rates in Section˜4 imply the trade-off between small and large values of the size $M$. If the size $M$ is large, the first term in the convergence rates—regardless of the deterministic or probabilistic one—is suppressed, while the second term amplifies. Conversely, if the size $M$ is small, the first term is not suppressed, while the second term does not amplify. In the extreme case where $M$ is equal to one, BIS is identical to standard self-normalized importance sampling by construction. Choosing a reasonable value of the size $M$ controls the trade-off, leading to an efficient approximation performance of BIS.
This subsection performs a simulation study on sensitivity of BIS to the size of the proposal sequence $M$. We used the experimental setting in Section˜6.1, while changing the size of the proposal sequence $M$. Figure˜S5 shows the approximation error of BIS for each benchmark density, under different values of the proposal-sequence size $M$. As anticipated, when the size $M$ was the small value $2$, 88, and $32$, the approximation performance of BIS got similar to that of standard importance sampling based on the scaled Halton sequence. In addition, the excessively large value of the proposal-sequence size $M=32768$ didn’t result in the best approximation performance of BIS. Overall, the size $M=2048$ and $M=8196$ led to the fastest decay of the approximation error of BIS. As the smaller size $M$ reduces the computation cost in BIS, we recommend the size $M=2048$ for experiments in two-dimensional domains based on this simulation study.
@@
Figure S5: The approximation error of BIS for each benchmark density when the size of the proposal sequence is $2$, 88, $32$, $128$, $512$, $2048$, $8196$, and $32768$. The experiment was repeated 10 times, where the bold line represents the averaged error and the band represents the 95% confidence interval.
@@ Appendix D Additional Experimental Details
This section provides additional details of the experiments presented in the main text. Section˜D.1 contains that of Section˜6.1 and Section˜D.2 contains that of Section˜6.2.
D.1 Additional Details of Section˜6.1
All the benchmark densities are given in the form proposed in Jarvenpaa2021 s.t.
[…]
@@
$k(\theta,\theta^{\prime})=\sigma^{2}\exp\left(-\frac{\|\theta-\theta^{\prime}\|^{2}}{2l^{2}}\right)$ (S85)
@@
$\operatorname{MMD}(p,q)^{2}=\mathbb{E}_{\theta,\theta^{\prime}\sim p}[\kappa(\theta,\theta^{\prime})]-2\mathbb{E}_{\theta\sim p,\theta^{\prime}\sim q}[\kappa(\theta,\theta^{\prime})]+\mathbb{E}_{\theta,\theta^{\prime}\sim q}[\kappa(\theta,\theta^{\prime})].$ (S86)
This alternative form of MMD can be efficiently estimated by Monte Carlo integration using samples from the densities $p$ and $q$. Throughout the experiments, we used MMD associated with the Gaussian kernel $\kappa(\theta,\theta^{\prime})=\exp(-0.5\|\theta-\theta^{\prime}\|^{2}/h)$ with the constant $h=0.1$. The expectation in the MMD with respect to each target density was approximated using sufficiently many 100,000 weighted samples drawn from the scaled Halton sequence via standard importance sampling.
@@
$\operatorname{TVD}(p,q):=\frac{1}{2}\int_{\Theta}\left|p(\theta)-q(\theta)\right|d\theta.$ (S87)
We approximate the integral over $\Theta$ by numerical integration. In this experiment, the densities $p$ and $q$ to be plugged in TVD are available only up to the normalizing constants. We also approximate their normalizing constants by numerical integration. Let $\{\eta_{i}\}_{i=1}^{N}$ be the $N$ points from the scaled Halton sequence on the domain $\Theta$. Denote the volume of the domain $\Theta$ by $V$. Let $\tilde{p}$ and $\tilde{q}$ be the proportional terms of the densities $p$ and $q$. Then their normalizing constants, denoted by $Z_{p}$ and $Z_{q}$, are approximated as follows:
$Z_{p}\approx\frac{V}{N}\sum_{i=1}^{N}\tilde{p}(\eta_{i})=:\hat{Z}_{p}\quad\text{and}\quad Z_{q}\approx\frac{V}{N}\sum_{i=1}^{N}\tilde{q}(\eta_{i})=:\hat{Z}_{q}.$ (S88)
@@
$\operatorname{TVD}(p,q)$ $\approx\frac{V}{2N}\sum_{i=1}^{N}\left|\frac{\tilde{p}(\eta_{i})}{\hat{Z}_{p}}-\frac{\tilde{q}(\eta_{i})}{\hat{Z}_{q}}\right|.$ (S89)
@@
Table˜S1 summarized the number of samples required for standard importance sampling to surpass the approximation error of 100 weighed samples obtained by BIS. For standard importance sampling, we used the scaled Halton sequence as samples and assigned self-normalized importance weights to them. It demonstrates that standard importance sampling required nearly 2,000 samples on average to achieve the approximation error that BIS does with 100 samples.
Table S1: The number of samples required for standard importance sampling to surpass the approximation error of 100 BIS weighed samples. The experiment was repeated 10 times, where the averaged number was reported together with the standard deviation in a bracket.
@@
D.2 Additional Details of Section˜6.2
The synthetic likelihood of the modified Lorentz weather model is computed using summary statistics of sample paths simulated from the model. Denote by $s\in\mathbb{R}^{d}$ summary statistics of the observed sample path, for some dimension $d$. Denote by $s(\theta)\in\mathbb{R}^{d}$ summary statistics of a sample path simulated from the model under a parameter $\theta$, where $s(\theta)$ is a random variable induced by simulation from the model. Let $m(\theta)\in\mathbb{R}^{d}$ and $\Sigma(\theta)\in\mathbb{R}^{d\times d}$ be the mean and covariance of $s(\theta)$. The synthetic likelihood at $\theta$ is then the Gaussian density $N(s\mid m(\theta),\Sigma(\theta))$ evaluated at the summary statistics $s$ of data. In practice, the mean and covariance are estimated from multiple i.i.d. realizations of the summary statistics $s(\theta)$. Thus, the synthetic likelihood is computed using multiple sample paths simulated from the model at each $\theta$. For the modified Lorentz weather model, Hakkarainen2012 used the following six summary statistics: (i) mean of $x_{k}(t)$ over time, (ii) variance of $x_{k}(t)$ over time, (iii) autocovariance of $x_{k}(t)$ with time lag one, (iv) covariance of $x_{k}(t)$ and $x_{k+1}(t)$ over time, (v) cross covariance of $x_{k}(t)$ and $x_{k-1}(t)$ with time lag one, (vi) cross covariance of $x_{k}(t)$ and $x_{k+1}(t)$ with time lag one. The summary statistics computed for each variable $x_{k}(t)$ is averaged over $k$. Thus, we have six-dimensional summary statistics for each sample path of the 40 variables.